OpenStudy (danielajohana):

solve the equation 2sin^2 (x)cos(x)=cos(x) on the interval [0,2pi)

OpenStudy (danielajohana):

\[2\sin ^{2}(x)\cos(x)=\cos(x)\]

Nnesha (nnesha):

Set it equal to zero and then take out the common factor

OpenStudy (danielajohana):

\[2\sin ^{2}(x)\cos(x)-\cos(x)=0\] like this?

Nnesha (nnesha):

yes right :)

Nnesha (nnesha):

so what is the common factor ?

OpenStudy (danielajohana):

At the beginning, couldn't I have just divided both sides by cos(x)? that would have gotten rid of it on both sides and have set it equal to zero @Nnesha

Nnesha (nnesha):

right!!! i don't why i was thinking that there is a sign between cos and sin you're right!

Nnesha (nnesha):

\[2\sin^2=0\] now solve for sin

Nnesha (nnesha):

wait hmm

Nnesha (nnesha):

hmm i don't think so... cuz if we divide both sides by cos(x)we will get just sin function

Nnesha (nnesha):

i would not divide by cos.

Nnesha (nnesha):

\[\huge\rm 2\sin^2(x)\cos(x)=\cos(x)\] is this your question ?

OpenStudy (danielajohana):

@Nnesha yes

Nnesha (nnesha):

yeah okay let's say sin=x and cos = y \[\huge\rm 2x^2y=y\] you would subtract y both sides not divide

OpenStudy (danielajohana):

yes, but since the question is asking to solve the equation, isn't it asking for when the equations intercept?

Nnesha (nnesha):

ye they are asking for solutions (x,y) <-- exact value of cos and sin between 0 to 2pi

Nnesha (nnesha):

\[\huge\rm \frac{ 2\sin^2(x)\cancel{\cos(x) }}{ \cancel{\cos(x}) } = \frac{ \cancel{\cos(x) }}{ \cancel{\cos(x)}}\] \[2\sin^2(x)=1\] but if you subtract cos(x) both sides you will get \[\large\rm \color{Red}{2\sin^2\cos(x)- \cos(x)=0}\]\[\cos(x)[2\sin^2(x)]=0\] so after that you should set both function equal to zero in this way you will get cos and sin both function value

Nnesha (nnesha):

well you confused me lol let me confirm

Nnesha (nnesha):

btw i'll go with 2nd method 1) subtract 2) find common factor :)

Nnesha (nnesha):

o.O

OpenStudy (danielajohana):

okay I understood how to do it from the comment before it was deleted

OpenStudy (freckles):

we can't divide both sides by 0 and cos(x) can be 0 -- but yeah you will end up with two equations to solve in the end: cos(x)=0 and 2 sin^2(x)=1

Nnesha (nnesha):

ahh freckles you here i was looking for u but you wer offline ;~;

OpenStudy (freckles):

\[2 \sin^2(x)\cos(x)-\cos(x)=0 \\ \cos(x)(2\sin^2(x)-1)=0 \\ \cos(x)=0 \text{ or } 2 \sin^2(x)-1=0 \\ \cos(x)=0 \text{ or } 2\sin^2(x)=1\]

Nnesha (nnesha):

forgot 1 there facepalm

OpenStudy (danielajohana):

okay, so for cos(x)=0 that would be pi/2 and 3pi/2 correct?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm \frac{ 2\sin^2(x)\cancel{\cos(x) }}{ \cancel{\cos(x}) } = \frac{ \cancel{\cos(x) }}{ \cancel{\cos(x)}}\] \[2\sin^2(x)=1\] but if you subtract cos(x) both sides you will get \[\large\rm \color{Red}{2\sin^2\cos(x)- \cos(x)=0}\]\[\cos(x)[2\sin^2(x)]=0\] so after that you should set both function equal to zero in this way you will get cos and sin both function value \(\color{blue}{\text{End of Quote}}\) correction \[\large\rm \color{Red}{2\sin^2\cos(x)- \cos(x)=0}\]\[\cos(x)[2\sin^2(x)-1]=0\]

Nnesha (nnesha):

ye that's right

OpenStudy (danielajohana):

yes, so the other one would be \[\sin ^{2}(x)=\frac{ 1 }{ 2 }\]

Nnesha (nnesha):

yep but that's sin^2 so take square root both sides

OpenStudy (danielajohana):

\[\sin(x)=\pm \frac{ \sqrt{2} }{ 2 }\] ?

Nnesha (nnesha):

yes right

OpenStudy (danielajohana):

okay so that's pi/4, 3pi/4, 5pi/4, and 7pi/4

Nnesha (nnesha):

nice

OpenStudy (danielajohana):

I get it thank you!

Nnesha (nnesha):

np :)